OpenStudy (anonymous):

SOMEONE PLEASE HELP I WILL FAN AND MEDAL!! Sin x = 2/5

OpenStudy (anonymous):

you want the angle whose sin has a value =2/5 so look inverse sin on calculator =23.57 deg.

OpenStudy (anonymous):

So the answer is 22.57?

OpenStudy (anonymous):

I'm solving the measurement of a right triangle with a hypotenuse of 5 and a opposite of 2 so the equation would be Sin X= 2/5 correct?? @jamaica25

OpenStudy (anonymous):

@Data_LG2 @mathmate can either of you help please!!!

OpenStudy (anonymous):

yes that's right : sin=opposite/hypotenuse get the inverse sin of both sides and you'll get x=23.58 degrees ^_^

OpenStudy (anonymous):

THANK YOU SO MUCH !!!!

OpenStudy (anonymous):

@Data_LG2 could you help me with one more please

OpenStudy (anonymous):

sure, i'll try :D

OpenStudy (anonymous):

Tan X= 2/1 Basically this is the same question as before just with tan instead of sin

OpenStudy (mathmate):

 Since tan(theta)=2/1 to find the angle, you look up \(\rm tan^{-1}(\frac{2}{1})\) It should be more than 45 degrees because tan(45 degrees) = 1.

OpenStudy (anonymous):

thank you so much you two are life savers

OpenStudy (mathmate):

You're welcome! :)